### v1

# MACE-Dance: Motion-Appearance Cascaded Experts for Music-Driven Dance Video Generation

Abstract
With the rise of online dance‑video platforms and rapid advances in AI‑generated content (AIGC), music‑driven dance generation task has emerged as a compelling research direction. Despite substantial progress in related domains such as music-driven 3D dance generation, pose-driven image animation, and audio-driven talking-head synthesis, they cannot be directly adapted to this task. Moreover, the limited studies in this task still struggle to jointly achieve high-quality visual appearance and human motion. Accordingly, we present MACE-Dance, a music-driven dance video generation framework with cascaded Mixture-of-Experts (MoE). The Motion Expert performs music-to-3D motion enforcing kinematic plausibility and artistic expressiveness, while the Appearance Expert carries out motion-and-reference conditioned video synthesis, preserving visual identity with spatiotemporal coherence. Specifically, the Motion Expert adopts Diffusion Model with BiMamba-Transformer hybrid architecture and Guidance-Free Training (GFT) strategy, achieving state-of-the-art (SOTA) performance in 3D dance generation task; the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy, achieving state-of-the-art (SOTA) performance in the pose-driven image animation task. To better benchmark this task, we curate a large-scale and diverse dataset, and design a motion–appearance evaluation protocol. Based on them, MACE-Dance also achieves the state-of-the-art (SOTA) performance. Project page: https://macedance.github.io/.
Figure 1: Leveraging the synergistic collaboration among the cascaded experts, MACE-Dance can generate diverse dance videos that not only exhibit kinematically plausible and artistically expressive motion, but also maintain spatiotemporal coherent appearance.
1 Introduction
Dance is a vital part of human culture. Moving to the beat and melody, dancers both convey emotion and narrative intent and showcase the power and beauty of human movement [37, 2, 1]. In the era of the internet, dance videos have become highly prominent on platforms such as YouTube and TikTok. In parallel, rapid advances [46, 48] in AI-generated content (AIGC) has created the technical preconditions for automating dance video creation, making it a timely and impactful research direction. Nevertheless, this task faces two key challenges [45]: (1) generating dance motions that are kinematically plausible while artistically expressive; and (2) achieving high-fidelity visual appearance with strong spatiotemporal consistency.
Recent progress in dance generation has focused primarily on 3D dance [37, 19, 17], with numerous strong methods emerging across model families—autoregressive [30, 47, 46], GAN-based [44, 35, 13], and diffusion-based [37, 19, 18]. Although 2D dance videos can be rendered from 3D motion, the resulting outputs typically miss realistic human–scene interactions and high-fidelity human textures, leading to a suboptimal visual appearance [45]. In contrast, human-centric image animation leverages a reference image along with various driving signals to generate videos that exhibit high-quality visual appearance. Traditionally, pose-driven image animation has achieved notable advances [36, 12, 5]. However, its utility for dance video generation is limited, as pose design—widely regarded as the most challenging and time-consuming step—remains manual [2]. Similarly, audio-driven talking head generation has also achieved significant breakthroughs [26, 27, 28]. However, its direct transfer to dance video generation remains challenging, as it primarily focuses on relatively simple upper-body gesture rather than the complex full-body motion required in dance [25, 24]. Research on music-driven dance video generation remains limited [4, 40], and there is still insufficient exploration of jointly achieving high-quality visual appearance and human motion.
Accordingly, we present MACE-Dance, a music-driven dance video generation framework with cascaded mixture-of-experts (MoE), as shown in Fig.   MACE-Dance: Motion-Appearance Cascaded Experts for Music-Driven Dance Video Generation . The Motion Expert performs music-to-3D motion enforcing kinematic plausibility and artistic expressiveness, while the Appearance Expert carries out motion-and-reference conditioned video synthesis, preserving visual identity with spatiotemporal coherence. (1) Motion Expert. Motion Expert adopts Diffusion Model with BiMamba-Transfomer hybrid architecture. The bidirectional Mamba [8] captures intra-modal local dependencies in music or dance, while the Transformer [39] models cross-modal global context. Owing to this architecture, the Motion Expert generates entire sequence in non-autoregressive manner during inference, not only improving generation efficiency, but also avoiding the exposure bias problem in autoregressive [47] and inpainting-based [37] methods. To enhance generation stability and accelerate inference, we employ guidance-free training (GFT [3]) instead of conventional classifier-free guidance (CFG [9]), enhancing the physical plausibility and artistic expressiveness for the generated dance. (2) Appearance Expert. Wan-Animate [5] has recently garnered substantial attention in both industry and academia. However, directly applying it to dance video generation yields limited effectiveness, as dance videos exhibit significantly more complex patterns than general videos. Thus, the Appearance Expert adopts a decoupled Kinematic–Aesthetic two-stage fine-tuning strategy to achieve high-fidelity appearance synthesis for dance videos. In Kinematic stage, it fine-tunes the Body Adapter to strengthen kinematic conditioning and motion adherence. In Aesthetic stage, it attaches a LoRA [11] branch to each DiT block and fine-tunes for aesthetic refinement, enhancing texture fidelity and stylistic consistency.
To better benchmark music-driven dance video generation task, we curate a large-scale and diverse dataset and design a motion–appearance evaluation protocol. Firstly, we curate a large-scale dance video dataset, named MA-Data, comprising  70k clips of 5–10 seconds each (totaling 116 hours), spanning over 20 dance genres. The dataset consists of two complementary sources: (1) 3D-rendered data (motion-centric): Derived from FineDance [17]—the largest 3D dance dataset recorded by professional dancers—we render front-view videos and extract random 5–10 s segments via a sliding window, yielding  20k clips (28 h). This subset emphasizes dance motion rather than visual appearance. (2) In-the-wild internet data (appearance-centric): Collected from high-engagement videos on platforms such as TikTok and YouTube, using the same sliding-window strategy to obtain  50k 5-10 s clips (88 h). This subset emphasizes visual appearance, while motions are relatively unprofessional. Secondly, we introduce a motion–appearance evaluation protocol. For the motion dimension, we assess the fidelity, diversity, and synchronization [16, 17] from Human-Kinematics perspective based on the 2D keypoints extracted by ViTPose [41]. For the appearance dimension, we leverage VBench [14] to evaluate key visual factors critical to dance videos, including imaging quality, aesthetic quality, subject consistency, background consistency, motion smoothness, temporal flickering.
In conclusion, our contributions are as follows:
  • To better benchmark the music-driven dance video generation task, we curate a large-scale and diverse dataset named MA-Data, along with a motion–appearance evaluation protocol.
  • Based on them, we introduce MACE-Dance, a music-driven dance video generation framework with cascaded experts, achieving state-of-the-art (SOTA) performance.
  • The Motion Expert adopts Diffusion Model with BiMamba-Transformer hybrid architecture and Guidance-Free Training strategy, achieving state-of-the-art (SOTA) performance in both quality and efficiency on the FineDance dataset in 3D dance generation task.
  • The Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy, achieving state-of-the-art (SOTA) performance on the MA-Data dataset in the pose-driven image animation task.
2 Related Work
2.1 Music-Driven 3D Dance Generation
Music and dance are deeply intertwined, and recent progress in music-to-dance generation has largely centered on 3D motion. Broadly, existing methods fall into three families: GAN-based, autoregressive, and diffusion-based models. 1) GAN-based models. Generators synthesize motion from music while discriminators provide adversarial feedback. Examples include CoheDancers [44], DeepDance [35], and Choreography cGAN [13]. 2) Autoregressive models. These methods typically adopt a two-stage pipeline: curating choreographic units by VQ-VAE [38] or FSQ [23], followed by autoregressive modeling of music-conditioned distributions over these units. Works such as Bailando [30], Bailando++ [31], Duolando [32], and MEGADance [47] fall into this paradigm. 3) Diffusion-based models. These methods corrupt motion with noise and train denoising networks to iteratively recover sequences conditioned on music, enabling diverse and temporally coherent dances. Representative works include EDGE [37], FineNet [17], Lodge [19], Lodge++ [18], and GCDance [20]. Despite substantial progress, 3D dance generation largely focuses on motion generation and underemphasizes visual appearance—an essential aspect of dance as an art form. Although 2D dance videos can be rendered from 3D motion, the outputs typically lack realistic human–scene interactions and high-fidelity human textures.
2.2 Human-Centric Image Animation
In contrast, human-centric image animation leverages a reference image along with various driving signals to generate videos that exhibit high-quality visual appearance, making it a promising direction for dance video generation. Firstly, pose-driven image animation utilizes 2D keypoints to generate motion videos, achieving notable advances [36, 12, 5], including Animate-X [36], Animate Anyone [12] and Wan-Animate [5]. However, its utility for dance video generation is limited, as pose design—widely regarded as the most challenging and time-consuming step—remains manual [2]. Secondly, speech-driven image animation employs audio features to generate talking head videos, also achieving significant breakthroughs [26, 27, 28], such as SyncTalk [26], OmniSync [28] and Hallo2 [6]. However, its direct transfer to dance video generation remains challenging, as these methods primarily focus on relatively simple upper-body gestures rather than the complex full-body motion required in dance [25, 24]. Finally, research on music-driven dance video generation remains limited  [4, 40]. DabFusion [40] introduces an optical-flow–prediction-based Diffusion Model; however, the generated videos exhibit blurry foreground subjects and backgrounds, thereby degrading visual fidelity. X-Dancer  [4] first predicts 2D keypoints from music and then drives image animation with these keypoints. Although this yields modest improvements, it overlooks the inherently 3D nature of dance; relying on a 2D representation severely constrains the model’s understanding of dancer locomotion, ultimately compromising motion quality. In conclusion, existing works for dance video generation still fail to simultaneously deliver high-quality visual appearance and human motion. Thus, leveraging synergistic collaboration among cascaded experts, MACE-Dance generates dance videos that not only exhibit kinematically plausible and artistically expressive motion but also maintain spatiotemporally coherent appearance.
3 Methodology
3.1 Overview
Given a music $M\in R^{T\times C_{m}}$ and reference image $I\in R^{H\times W\times 3}$, our objective is to synthesize the corresponding dance videos $D\in R^{T\times H\times W\times 3}$ with high-quality visual appearance and human motion. Overall, MACE-Dance is a music-driven dance video generation framework with cascaded mixture-of-experts (MoE), as shown in Fig.  2. The Motion Expert (ME) transfers music sequence $M$ into 3D motion sequence $X\in R^{T\times C_{x}}$, enforcing kinematic plausibility and artistic expressiveness. The Appearance Expert (AE) utilizes the above 3D motion sequence $X$ and reference image $I$ to drive video synthesis, preserving visual identity with spatiotemporal coherence.
Figure 2: Overview of MACE-Dance.Leveraging the cascaded Mixture-of-Experts (MoE) design, the Motion Expert generates kinematically plausible and artistically expressive 3D motion $X$ conditioned on the music $M$, and the Appearance Expert then animates the reference image II with the 3D motion $X$, yielding the dance video $D$ that exhibits spatiotemporally coherent appearance. Thanks to the Guidance-Free Training (GFT) strategy, $\beta\in[0,1]$ can serve as a controllable knob that governs the diversity of the generated motion.
3.2 Motion Expert
3.2.1 Generative Strategy.
Diffusion. DDPM [10] defines diffusion as a Markov noising process with latents $\{z_{t}\}_{t=0}^{T}$ that follow a forward noising process $q(z_{t}|x)$, where $x\sim p(x)$ is drawn from the 3D dance data distribution. The forward noising process is defined as:
$q(z_{t}|x)\sim\mathcal{N}(\sqrt{\bar{\alpha}_{t}}x,(1-\bar{\alpha}_{t})I),$ (1)
where $\bar{\alpha}_{t}\in(0,1)$ are constants which follow a monotonically decreasing schedule such that when $\bar{\alpha}_{t}$ approaches 0. Timestep $T$ are commonly set to 1000, and $z_{T}\sim\mathcal{N}(0,I)$. With paired music conditioning $c$, we can reverse the forward diffusion process by learning to estimate $\hat{x}_{\theta}(z_{t},t,c)\approx x$ with model parameters $\theta$ for all $t$. We can optimize $\theta$ by the naive reconstruction loss in Diffusion Model [10]:
$\mathcal{L}_{\text{DM}}=\mathbb{E}\big[\|\hat{x}_{\theta}(z_{t},t,c)-x\|_{2}^{2}\big].$ (2)
Guidance-Free Training. While conventional Classifier-Free Guidance (CFG [9]) provides a knob for trading off diversity against fidelity, it is notoriously sensitive to the guidance scale. Suboptimal settings can easily disrupt this balance, manifesting in 3D motion as excessive jitter or repetitive, monotonous movements, ultimately degrading the dance quality. Moreover, the search for a reliable scale also incurs substantial tuning overhead. Motivated by these limitations, we adopt Guidance-Free Training (GFT [3]). GFT reformulates conditional training to directly learn a $\beta$-indexed sampling model via linear interpolation with the unconditional output. This allows a single model to capture an entire family of diversity-fidelity trade-offs robustly, eliminating the need for post-hoc guidance. Accordingly, we establish $x_{\beta}$ as the new optimization target for our model $\theta$:
$x_{\beta}=\beta\hat{x}_{\theta}(z_{t},t,c,\beta)+(1-\beta)\mathbf{sg}[\hat{x}_{\theta}(z_{t},t,\emptyset,1)]$ (3)
where $\emptyset$ denotes the unconditional setting, and $\mathbf{sg}$ represents the stop-gradient operation. $\beta$ serves as a temperature parameter that is also provided to the model $\theta$ as an additional conditioning input. During training, $\beta$ and $t$ are sampled randomly from $U(0,1)$ and the integer set $\{0,1,\dots,T\}$, respectively. Moreover, we further apply the reconstruction loss, 3D joint loss, velocity loss, foot contact loss, to enhance physical plausibility and aesthetic expressiveness:
$\mathcal{L}_{\text{rec}}$ $=\mathbb{E}\big[\|x_{\beta}-x\|_{2}^{2}\big],$ (4)
$\mathcal{L}_{\text{joint}}$ $=\mathbb{E}\big[\|FK(x_{\beta})-FK(x)\|_{2}^{2}\big],$
$\mathcal{L}_{\text{vel}}$ $=\mathbb{E}\big[\|FK(x_{\beta})^{\prime}-FK(x)^{\prime}\|_{2}^{2}\big]$
$\mathcal{L}_{\text{foot}}$ $=\mathbb{E}\big[\|FK(x_{\beta})^{\prime}\cdot\hat{\mathbf{b}}\|_{2}^{2}\big],$
where $FK(\cdot)$ denotes the forward kinematic function that converts joint angles into joint positions, and $\hat{\mathbf{b}}$ is the model’s own prediction of the binary foot contact label’s portion of the pose. Our overall training loss $\mathcal{L}$ is the weighted sum of the above losses, where the weights $\lambda$ were chosen to balance the magnitudes of the losses:
$\mathcal{L}=\lambda_{\text{rec}}\mathcal{L}_{\text{rec}}+\lambda_{\text{joint}}\mathcal{L}_{\text{joint}}+\lambda_{\text{vel}}\mathcal{L}_{\text{vel}}+\lambda_{\text{foot}}\mathcal{L}_{\text{foot}}.$ (5)
Inference. At each of the denoising timesteps $t$, Motion Expert predicts the denoised sample and noises it back to timestep t1t-1: $\hat{z}_{t-1}\sim q(\hat{x}_{\theta}(z_{t},t,c,\beta),t-1)$, terminating when it reaches $t=0$. We utilize Denoising Diffusion Implicit Models (DDIM [33]) to accelerate the sampling procedure. Values of $\beta$ near 0 favor high fidelity, while values near 1 favor high diversity. Thus, $\beta$ can also be regarded as a control signal, and we set its value to 0.75. Notably, GFT inherently achieves nearly double the generation efficiency compared to conventional CFG, as it only requires a single conditional computation per step, eliminating the need for simultaneous conditional and unconditional predictions.
3.2.2 Model Architecture
Overview. Motion Expert adopts a BiMamba–Transformer hybrid backbone, thereby enabling the generation of temporally coherent and musically aligned dance motions. BiMamba captures intra-modal local dependencies in music or dance, while the Transformer models cross-modal global context. As shown in Fig.  2, the architecture details are as follows: Firstly, our model conditions the generator on the Librosa [22]-extracted music features from $M$ as [16], which are then processed by an $L_{m}$‑layer BiMamba to capture intra‑modal temporal dynamics. Secondly, the diffusion time step $t$ and temperature parameter $\beta$ are encoded as sinusoidal embeddings and fused by element-wise addition to yield a $t$-$\beta$ embedding used throughout the generator. Third, the dance generator consists of $L_{d}$ stacked blocks. In each block: (1) the current state $z_{t}$ is first passed through a BiMamba to model intra-modal local dependencies; (2) FiLM  [29] is applied to modulate the features with the fused $t$-$\beta$ embedding; (3) a Transformer performs cross-modal attention over the music encoding to integrate global musical context, and subsequently passes the result through a feed-forward network; and (4) a second FiLM [29] further reinforces the $t$-$\beta$ conditioning. Finally, the generator outputs the 3D motion sequence $\hat{x}_{\theta}(z_{t},t,c,\beta)$ (i.e. $X$), represented as SMPL [21] parameters. Owing to this architecture, the Motion Expert generates the entire sequence in a non-autoregressive manner during inference, not only improving generation efficiency but also avoiding the exposure-bias problem in autoregressive [47] and inpainting-based [37] methods.
Intra-Modal Local-Dependency. While the Transformer excels at temporal modeling, it is inherently position-invariant and captures sequence order only through positional encodings [39], which limits its deep understanding of local dependencies. In contrast, music-to-dance generation demands strong local continuity between movements. Owing to its inherent sequential inductive bias, Mamba [8] has demonstrated strong performance in modeling fine-grained local dependencies  [42, 7]. Moreover, Bidirectional Mamba processes inputs in both forward and backward directions, enabling wider representations and deeper understanding of music and dance. Specifically, the Selective State Space Model (Mamba) integrates a selection mechanism and a scan module (S6) [8] to dynamically emphasize salient input segments for efficient sequence modeling. Unlike traditional SSMs with time-invariant parameters, Mamba generates input-dependent $\bar{A}_{t},\bar{B}_{t},C_{t}$ through fully connected layers, enhancing generalization. For each time step $t$, the input $x_{t}$, hidden state $h_{t}$, and output $y_{t}$ evolve as:
$h_{t}=\bar{A}_{t}h_{t-1}+\bar{B}_{t}x_{t},\quad y_{t}=C_{t}h_{t},$ (6)
where $\bar{A}_{t},\bar{B}_{t},C_{t}$ are dynamically updated, and the state transitions become:
$\bar{A}=\exp(\Delta A),\quad\bar{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B,$ (7)
where $\Delta$ is the discretization step size, $A$ is the continuous-time state transition matrix, $B$ is the input projection matrix, and $C$ is the output projection matrix.
Cross-Modal Global-Context. While BiMamba [8] excels at capturing local dependencies, we employ a Transformer [39] block to model the cross-modal global context, which is crucial for aligning the overall dance structure with long-term musical phrasing. This block consists of a cross-attention layer followed by a feed-forward network (FFN). The attention layer is the core component as:
$\text{Attention}=\text{softmax}\left(\frac{Q_{d}\cdot K_{m}^{T}}{\sqrt{C}}\right)V_{m}.$ (8)
Table 1: Quantitative comparison with SOTAs on the MA-Data dataset in Music-Driven Dance Video Generation task.
Appearance Motion
[…]
Ground Truth 67.12 53.51 91.86 92.97 98.20 96.88 9.24 5.31 0.526
Hallo2 [6] [ICLR’25] 62.64 50.79 92.48 93.84 98.30 96.56 16.55 1.29 8.11 5.47 0.505
EDGE [17] [CVPR’23] 63.05 49.70 91.79 93.30 98.64 97.10 21.77 1.39 9.08 5.74 0.498
Lodge [19] [CVPR’24] 63.69 49.22 91.67 92.98 98.46 97.05 18.73 1.49 8.87 5.71 0.474
MEGA [47] [NeurIPS’25] 66.14 49.89 92.95 94.13 97.45 96.32 18.98 1.65 8.78 5.59 0.513
MACE-Dance 65.35 51.79 93.97 94.57 98.46 97.10 16.46 0.28 9.74 6.34 0.523
Table 2: Quantitative comparison with SOTAs on the FineDance dataset in Music-Driven 3D Dance Generation task.
[…]
Ground Truth 0.216 9.94 7.54 0.201
FACT [16] [ICCV’21] 113.38 97.05 0.284 3.36 6.37 0.183 29
MNET [15] [CVPR’22] 104.71 90.31 0.394 3.12 6.14 0.186 26
Bailando [30] [CVPR’22] 82.81 28.17 0.188 7.74 6.25 0.202 188
EDGE [37] [CVPR’23] 94.34 50.38 0.200 8.13 6.45 0.212 119
Lodge [19] [CVPR’24] 50.00 35.52 0.028 5.67 4.96 0.226 224
MEGA [47] [NeurIPS’25] 50.00 13.02 0.243 6.23 6.27 0.226 238
GFT $\rightarrow$ CFG 25.54 36.31 0.295 11.57 7.03 0.223 475
BiMamba $\rightarrow$ Mamba 65.10 51.74 0.35 9.59 7.66 0.224 1044
BiMamba $\rightarrow$ Transformer 104.93 114.42 0.12 7.90 4.22 0.266 683
Motion Expert (Full) 17.83 25.09 0.210 10.30 8.09 0.229 770
3.3 Appearance Expert
Wan-Animate [5] has recently garnered substantial attention in both industry and academia. However, it is designed for general-purpose motion synthesis; direct transfer to dance video generation is suboptimal due to the domain gap and the richer spatiotemporal complexity of dance—namely intricate whole-body coordination and dynamic camera choreography. Accordingly, the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy to achieve high-fidelity appearance synthesis for dance videos.
Model Architecture. As illustrated in Fig. 2, the architecture of our Appearance Expert is built upon the Wan-Animate [5], which takes a reference image $I$ for appearance and a 3D motion sequence $X$ for motion guidance. The motion sequence $X$ is first projected to 2D keypoints, which are then encoded by a Body Adapter to yield motion features. These features are subsequently fused with the latent extracted from the reference image $I$. The resulting latent is processed by a backbone of stacked DiT blocks, where lightweight LoRA adapters are integrated into each block. The facial processing pipeline remains identical to that of Wan-Animate and is therefore omitted for clarity.
Kinematic Stage. In dance, body pose is paramount. The original Wan‑Animate prioritizes facial cues, allocating a dedicated cross‑attention branch to the face while fusing body signals only via additive injection. We therefore strengthen kinematic conditioning by fine‑tuning the Body Adapter in the Kinematic Stage to reweight and calibrate body features across scales, thereby enforcing motion adherence without altering the backbone. We intentionally do not introduce an additional body cross‑attention branch because (i) it disturbs the pretrained inductive bias and can compete with the facial cross‑attention, causing feature entanglement, (ii) it adds substantial memory/latency overhead and training instability on long, fast dance sequences.
Aesthetic Stage. To refine visual quality without disturbing motion control, we freeze the kinematic pathways and attach lightweight LoRA adapters to the attention (query/key/value/output) and feed-forward projections in each DiT block of Wan‑Animate. These rank‑r adapters enable parameter‑efficient specialization toward dance‑specific aesthetics—sharpening textures (skin, hair, fabric), stabilizing clothing and accessories, and handling rich camera choreography (pans, zooms, handheld motion)—while preserving pretrained content priors. Specifically, LoRA is an effective technique for adapting large pre-trained models to down-streaming tasks with few training-able parameters. To achieve this goal, LoRA introduces a low-rank decomposition-based method to the model’s weight matrix, enabling efficient adaptation to new tasks while maintaining the model’s original capabilities. Given the weight matrix $W_{0}\in\mathbb{R}^{m\times n}$ of the original pre-trained model, LoRA [11] uses two low-rank matrices $A\in\mathbb{R}^{m\times r}$ ($r\ll m$) and $B\in\mathbb{R}^{r\times n}$ ($r\ll n$) to shift the trained distribution according to the new data training. Thanks to the low-rank matrix in $A$ and $B$, LoRA updates the model more efficiently than the full rank matrix and shows comparable results with full training. Formally, the new weight matrix $W$ can be represented as:
$W=W_{0}+\Delta W=W_{0}+AB.$ (9)
4 Experiment
4.1 Dataset
To support music-driven dance video generation task, we curate a large-scale dance video dataset, named MA-Data. It comprises  70k clips of 5–10 seconds each (totaling 116 hours), and spans over 20 distinct dance genres, such as Jazz, Latin, Eastern Folk. MA-Data consists of two complementary sources. (1) 3D-rendered data (motion-centric). This subset is derived from FineDance [17], the largest 3D dance dataset recorded by professional dancers, and emphasizes dance motion rather than visual appearance. Specifically, we obtain front-view videos from 3D motion by 3D rendering technique, and extract random 5–10 s segments via a sliding window for data augmentation, yielding  20k clips ( 28 h). (2) In-the-wild internet data (appearance-centric). This subset is curated from high-engagement creators on platforms such as TikTok and YouTube, emphasizing visual appearance; motions typically prioritize entertainment value over technical rigor. As raw crawls include many samples misaligned with our task, we apply a multi-stage cleaning pipeline: (i) perform shot boundary detection with TransNet V2 [34], segment accordingly, and discard segments shorter than 5 s; (ii) remove near-static clips using an optical-flow magnitude threshold; (iii) enforce a single-performer constraint via ViTPose [41] by dropping clips containing two or more people; and (iv) split long videos into 5–10 s clips with a sliding window and random offsets. The final set comprises  50k clips ( 88 hours).
Table 3: Quantitative comparison with SOTAs on the MA-Data dataset in Pose-Driven Image Animation task.
[…]
Animate-Anyone [12] [CVPR’24] 515.26 0.648 0.091 19.65
Magic-Animate [43] [CVPR’24] 1032.06 0.311 0.207 14.00
Wan-Animate [5] [Arxiv’25] 332.82 0.707 0.078 21.11
w/o. Kinematic stage 328.91 0.596 0.107 18.69
w/o. Aesthetic stage 445.93 0.563 0.121 17.89
Appearance Expert 274.94 0.739 0.066 22.40
Figure 3: Qualitative comparison with SOTAs across reference image domains (real-person vs. anime-character) and music genres (Eastern Folk vs. Popping) in the music-driven dance video generation task.
4.2 Evaluation
The key challenges of music-driven dance video generation are: (1) generating dance motions that are kinematically plausible while artistically expressive; and (2) achieving high-fidelity visual appearance with strong spatiotemporal consistency. Inspired by this, we introduce a motion–appearance evaluation protocol. (1) Motion dimension. We first extract 2D keypoint sequences using ViTPose [41] from the dance videos and evaluate from a Human-Kinematics perspective. To evaluate the fidelity and diversity, we report FID and DIV across two feature spaces [16, 30]: (1) kinetic (k), capturing motion dynamics, and (2) geometric (g), encoding spatial joint relations. To measure music–motion synchronization, we utilize the Beat Alignment Score (BAS) [16, 19]. (2) Appearance dimension. We adopt a set of metrics from VBench [14], specifically choosing those most critical for the nuances of dance videos. Our evaluation includes imaging quality (IQ), aesthetic quality (AQ), subject consistency (SC), background consistency (BC), motion smoothness (MS), temporal flickering (TF). Finally, we collect an additional 500 5-second clips to construct the test set, with high engagement on TikTok and across multiple dance genres.
4.3 Comparison
Music-Driven Dance Video Generation. As there is currently no open-source implementation for Music-Driven Dance Video Generation, we compare MACE-Dance against two baseline families on the MA-Data dataset: (1) 3D dance generation methods pipelined with Wan-Animate, including EDGE [37], Lodge [19], and MEGA [47]; (2) talking-head generation methods adapted by replacing face masks with full-body masks followed by fine-tuning on the MA-Data dataset, exemplified by Hallo2 [6]. As shown in Tab. 1, our proposed MACE-Dance demonstrates state-of-the-art (SOTA) performance in both appearance and motion quality. Specifically, for the motion aspect, MACE-Dance achieves the best results across all metrics ($FIK_{k}=16.46,FID_{g}=0.28,DIV_{k}=9.74,DIV_{g}=6.34,BAS=0.523$); for the appearance aspect, it also attains best performance on most metrics, with scores of IQ=65.35,AQ=51.79,SC=93.97,BC=94.57,MS=98.46,IQ=65.35,AQ=51.79,SC=93.97,BC=94.57,MS=98.46, and $TF=97.10$. By effectively decoupling the task into 3D Dance Generation and Pose-Driven Image Animation, and leveraging the strong performance of the Motion Expert and Appearance Expert on their respective sub-tasks, MACE-Dance delivers exceptional generation quality. Note: User Study can be found in the supplementary material.
Music-Driven 3D Dance Generation. Music-Driven 3D Dance Generation is a canonical task and determines the motion quality of MACE-Dance. We compare the Motion Expert against FACT [16], MNET [15], Bailando [30], EDGE [37], Lodge [19], and MEGA [47] on the FineDance dataset with metrics following  [19]. As shown in Tab. 2, the Motion Expert attains overall state-of-the-art (SOTA) performance. Specifically, it achieves the best $FID_{k}=17.83$ and a competitive $FID_{g}=25.09$, indicating high fidelity; the best $DIV_{k}=10.30$ and $DIV_{g}=8.09$, indicating strong diversity; a competitive FSR, supporting physical plausibility; the best $BAS=0.229$, demonstrating superior audio-motion synchronization; and a substantially higher $FPS=770$, evidencing excellent generation efficiency. These advances primarily stem from: (1) adopting a Diffusion Model with a BiMamba-Transformer hybrid architecture, enabling high-quality long motion sequences in a non-autoregressive manner; and (2) a Guidance-Free Training (GFT) strategy that improves generation quality without requiring dual-pass inference (conditioned + unconditioned).
Pose-Driven Image Animation. Pose-Driven Image Animation is likewise a canonical task and governs the appearance quality of MACE-Dance. We compare the Appearance Expert against Animate-Anyone [12], Magic-Animate [43], and Wan-Animate [5] on the MA-Data dataset with metrics following  [5]. As shown in Tab. 3, the Appearance Expert achieves state-of-the-art (SOTA) performance, on all metrics FVD=274.94,SSIM=0.739,LPIPS=0.066,PSNR=22.40FVD=274.94,SSIM=0.739,LPIPS=0.066,PSNR=22.40. Its strong video synthesis quality primarily benefits from Wan-Animate (Baseline)’s powerful cross-modal understanding and our proposed Kinematic-Aesthetic decoupled fine-tuning strategy. Note: Qualitative analysis can be found in the supplementary material.
4.4 Qualitative Analysis
We also present a qualitative comparison against other methods across reference-image domains (real person vs. anime character) and music genres (elegant and rhythmically rich Eastern Folk vs. powerful and funk-inspired Popping) for the music-driven dance video generation task. As shown in Fig.  3, Hallo2 exhibits significant blurring in human details and noticeable artifacts in the background ; EDGE frequently produces abnormal movements that violate physical plausibility; Lodge often exhibits abrupt motion discontinuities, leading to poor visual appearance; and MEGA tends to generate overly simple and repetitive motions that lack artistic expressiveness. In contrast, videos generated by our method (MACE-Dance) not only present kinematically plausible and artistically expressive motion, but also maintain spatiotemporally coherent appearance. In summary, MACE-Dance demonstrates superior performance in qualitative evaluations.
Figure 4: Ablation for the model architecture of Motion Expert.
4.5 Ablation
[…]
Figure 5: Ablation for the Appearance Expert.
Appearance Expert Since the two-stage fine-tuning strategy is central to our Appearance Expert, we evaluate the contribution of each stage via ablation, as summarized in Tab.  3 and Fig. 5. (1) Kinematic Stage. We fine-tune the Body Adapter while freezing all other components. Removing this stage leads to a modest decline across all metrics quantitatively and noticeable kinematic errors and motion blur qualitatively, indicating its effectiveness for ensuring human kinematic plausibility in video. (2) Aesthetic Stage. We fine-tune LoRA parameters in each DiT block. Omitting this stage causes a substantial degradation across all metrics quantitatively and obvious ghosting artifact qualitatively, underscoring its critical role to preserve video aesthetic. Finally, Appearance Expert also demonstrates the superior performance over the Wan-Animate (Baseline), which validates the overall effectiveness of the proposed Kinematic-Aesthetic fine-tuning strategy.
5 Conclusion
In conclusion, we present MACE-Dance, a music-driven dance video generation framework with cascaded Mixture-of-Experts (MoE). The Motion Expert enforces kinematic plausibility and artistic expressiveness, while the Appearance Expert preserves visual identity with spatiotemporal coherence. Specifically, the Motion Expert adopts Diffusion Model with a BiMamba–Transformer hybrid backbone and Guidance-Free Training strategy, while the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy. To better benchmark this task, we curate a large-scale and diverse dataset, and desigm a motion–appearance evaluation protocol.Extensive experiments demonstrate the superiority of MACE-Dance and of its Motion and Appearance Experts. For future work, we plan to extend MACE-Dance with textual descriptions to enable more interactive and flexible dance generation, as text serves as the most flexible and intuitive form for specifying user needs.