### v3

# MACE-Dance: Motion-Appearance Cascaded Experts for Music-Driven Dance Video Generation

Abstract.
With the rise of online dance‑video platforms and rapid advances in AIGC, music‑driven dance generation task has emerged as a compelling research direction. Despite substantial progress in related domains such as music-driven 3D dance generation, pose-driven image animation, and audio-driven talking-head synthesis, these approaches are not readily transferable to this task due to fundamental mismatches in generation targets and constraints. Moreover, research on music-driven dance video generation remains limited and fails to capture the inherently 3D nature of dance, resulting in compromised motion quality and visual appearance. Accordingly, we present MACE-Dance, a music-driven dance video generation framework with cascaded Mixture-of-Experts (MoE). The Motion Expert performs music-to-3D motion enforcing kinematic plausibility and artistic expressiveness, while the Appearance Expert carries out motion-and-reference conditioned video synthesis, preserving visual identity with spatiotemporal coherence. Specifically, the Motion Expert adopts Diffusion Model with BiMamba-Transformer hybrid architecture and Guidance-Free Training (GFT) strategy, achieving state-of-the-art (SOTA) performance in 3D dance generation task; the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy, achieving state-of-the-art (SOTA) performance in the pose-driven image animation task. To better benchmark this task, we curate a large-scale dataset, and design a motion–appearance evaluation protocol. Based on them, MACE-Dance also achieves the state-of-the-art (SOTA) performance. Code is available at https://github.com/AMAP-ML/MACE-Dance.
copyright: noneccs: Applied computing Arts and humanitiesccs: Computing methodologies Computer visionccs: Human-centered computing
Figure 1. Leveraging the synergistic collaboration among the cascaded experts, MACE-Dance can generate diverse dance videos that not only exhibit kinematically plausible and artistically expressive motion, but also maintain spatiotemporal coherent appearance.
1. Introduction
Dance is a vital part of human culture. Moving to the beat and melody, dancers both convey emotion and narrative intent and showcase the power and beauty of human movement (Tseng et al., 2023; Butterworth*, 2004). In the era of the internet, dance videos have become highly prominent on platforms such as YouTube and TikTok. In parallel, rapid advances (Yang et al., 2025a; Zhuo et al., 2023; Chen et al., 2025a, b; Lei et al., 2025) in AI-generated content (AIGC) has created the technical preconditions for automating dance video creation, making it a timely and impactful research direction. Nevertheless, this task faces two key challenges: (1) generating dance motions that are kinematically plausible while artistically expressive; and (2) achieving high-fidelity visual appearance with strong spatiotemporal consistency.
Recent progress in dance generation has focused primarily on 3D dance (Tseng et al., 2023; Li et al., 2024c, 2023), with numerous strong methods emerging across model families-autoregressive (Siyao et al., 2022; Yang et al., 2025b, a), GAN-based (Yang et al., 2024b; Sun et al., 2019; Huang and Liu, 2021), and diffusion-based (Tseng et al., 2023; Li et al., 2024c, b). Although 2D dance videos can be rendered from 3D motion, such renderings typically lack realistic human–scene interactions and detailed appearance cues, resulting in visually suboptimal outputs (Yang et al., 2024c). In contrast, human-centric image animation leverages a reference image along with various driving signals to generate videos. Traditionally, pose-driven image animation has achieved notable advances (Tan et al., 2024; Hu, 2024; Cheng et al., 2025). However, its utility for dance video generation is limited, as pose design—widely regarded as the most challenging and time-consuming step—still remains manual (Butterworth*, 2004). Similarly, audio-driven talking head generation has also achieved significant breakthroughs (Peng et al., 2024, 2025b, 2025c). However, its direct transfer to dance video generation remains challenging, as it primarily focuses on relatively simple upper-body gesture rather than the complex full-body motion required in dance (Peng et al., 2023). Research on music-driven dance video generation remains limited (Chen et al., 2025e; Wang et al., 2025b; Tang et al., 2025) and fails to capture the inherently 3D nature of dance, resulting in compromised motion quality and visual appearance.
Accordingly, we present MACE-Dance, a music-driven dance video generation framework with cascaded mixture-of-experts (MoE), as shown in Fig.  1. The Motion Expert performs music-to-3D motion enforcing kinematic plausibility and artistic expressiveness, while the Appearance Expert carries out motion-and-reference conditioned video synthesis, preserving visual identity with spatiotemporal coherence. Notably, MACE-Dance adopts 3D SMPL (Loper et al., 2023) parameters rather than 2D keypoints as the intermediate representation, as 3D provides view-invariant and physically consistent supervision, while 2D projections introduce irreversible information loss and viewpoint ambiguity. (1) Motion Expert. Motion Expert adopts Diffusion Model with BiMamba-Transfomer hybrid architecture. The bidirectional Mamba (Gu and Dao, 2023) captures intra-modal local dependencies in music or dance, while the Transformer (Vaswani, 2017) models cross-modal global context. Owing to this architecture, the Motion Expert generates entire sequence in non-autoregressive manner during inference, not only improving generation efficiency, but also avoiding the exposure bias problem in autoregressive (Yang et al., 2025b) and inpainting-based (Tseng et al., 2023) methods. To enhance generation stability and accelerate inference, we employ guidance-free training (GFT (Chen et al., 2025c)) instead of conventional classifier-free guidance (CFG (Ho and Salimans, 2022)), enhancing the physical plausibility and artistic expressiveness for the generated dance. (2) Appearance Expert. Wan-Animate (Cheng et al., 2025) has recently garnered substantial attention in both industry and academia. However, directly applying it to dance video generation yields limited effectiveness, as dance videos exhibit significantly more complex patterns than general videos. Thus, the Appearance Expert adopts a decoupled Kinematic–Aesthetic two-stage fine-tuning strategy to achieve high-fidelity appearance synthesis. In Kinematic stage, it fine-tunes the Body Adapter to strengthen kinematic conditioning and motion adherence. In Aesthetic stage, it attaches a LoRA (Hu et al., 2022) branch to each DiT block and fine-tunes for aesthetic refinement, enhancing texture fidelity and stylistic consistency.
To better benchmark music-driven dance video generation task, we curate a large-scale dataset and design a motion–appearance evaluation protocol. Firstly, we curate a large-scale dance video dataset, named MA-Data, comprising  70k clips of 5–10 seconds each (totaling 116 hours), spanning over 20 dance genres. The dataset consists of two complementary sources: (1) 3D-rendered data (motion-centric): Derived from FineDance (Li et al., 2023)—the largest 3D dance dataset recorded by professional dancers—we render front-view videos and extract random 5–10 s segments via a sliding window, yielding  20k clips (28 h). This subset emphasizes professional dance motion rather than visual appearance. (2) In-the-wild internet data (appearance-centric): Collected from high-engagement videos on platforms such as TikTok and YouTube, using the same sliding-window strategy to obtain  50k 5-10 s clips (88 h). This subset emphasizes visual appearance, while motions are relatively unprofessional. Secondly, we introduce a motion–appearance evaluation protocol. For the motion dimension, we assess the fidelity, diversity, and synchronization (Li et al., 2021, 2023) from Human-Kinematics perspective based on the 2D keypoints extracted by ViTPose (Xu et al., 2022). For the appearance dimension, we adopt VBench (Huang et al., 2024)—a widely used benchmark in video generation—and select a set of dance-specific metrics.
In conclusion, our contributions are as follows: (1) To better benchmark the music-driven dance video generation task, we curate a large-scale dataset named MA-Data, along with a motion–appearance evaluation protocol. (2) Based on them, we introduce MACE-Dance, a music-driven dance video generation framework with cascaded experts, achieving SOTA performance. (3) The Motion Expert adopts Diffusion Model with BiMamba-Transformer hybrid architecture and Guidance-Free Training strategy, achieving SOTA performance on the FineDance dataset in music-driven 3D dance generation task. (4) Appearance Expert adopts a decoupled Kinematic-Aesthetic fine-tuning strategy, achieving SOTA performance on the MA-Data dataset in the pose-driven image animation task.
2. Related Work
2.1. Music-Driven 3D Dance Generation
Music and dance are deeply intertwined, and recent progress in music-to-dance generation has largely centered on 3D motion. Broadly, existing methods fall into three families: GAN-based, autoregressive, and diffusion-based models. 1) GAN-based models. Generators synthesize motion from music while discriminators provide adversarial feedback. Examples include CoheDancers (Yang et al., 2024b) and DeepDance (Sun et al., 2019). 2) Autoregressive models. These methods typically adopt a two-stage pipeline: curating choreographic units by VQ-VAE (van den Oord et al., 2017) or FSQ (Mentzer et al., 2023), followed by autoregressive modeling of music-conditioned distributions over these units (Yang et al., 2024a, 2026; Li et al., 2024a). Works such as Bailando (Siyao et al., 2022), Bailando++ (Siyao et al., 2023), and MEGADance (Yang et al., 2025b) fall into this paradigm. 3) Diffusion-based models. These methods corrupt motion with noise and train denoising networks to iteratively recover sequences conditioned on music (Yang et al., 2025c), enabling diverse and temporally coherent dances. Representative works include EDGE (Tseng et al., 2023), FineNet (Li et al., 2023), Lodge (Li et al., 2024c), Lodge++ (Li et al., 2024b), and GCDance (Liu et al., 2025). Despite substantial progress, 3D dance generation only focuses on motion generation and underemphasizes visual appearance—an essential aspect of dance as an art form. Although 2D dance videos can be rendered from 3D motion, the outputs typically lack realistic human–scene interactions and high-fidelity human textures.
2.2. Human-Centric Image Animation
In contrast, human-centric image animation leverages a reference image along with various driving signals to generate videos that exhibit high-quality visual appearance, making it a promising direction for dance video generation. Firstly, pose-driven image animation utilizes 2D keypoints to generate motion videos, achieving notable advances (Tan et al., 2024; Hu, 2024; Cheng et al., 2025), including Animate-X (Tan et al., 2024), Animate Anyone (Hu, 2024) and Wan-Animate (Cheng et al., 2025). However, its utility for dance video generation is limited, as pose design—widely regarded as the most challenging and time-consuming step— still remains manual (Butterworth*, 2004). Secondly, speech-driven image animation employs audio features to generate talking head videos, also achieving significant breakthroughs (Peng et al., 2024, 2025b, 2025c), such as SyncTalk (Peng et al., 2024), OmniSync (Peng et al., 2025c) and Hallo2 (Cui et al., 2024). However, its direct transfer to dance video generation remains challenging, as these methods primarily focus on relatively simple upper-body gestures rather than the complex full-body motion required in dance (Peng et al., 2023; Zhang et al., 2025c, b, d). Finally, research on music-driven dance video generation remains limited. DabFusion (Wang et al., 2025b) introduces an end-to-end Diffusion-based method, but the generated videos exhibit blurry foreground subjects and backgrounds, thereby degrading visual fidelity. X-Dancer  (Chen et al., 2025e), STG-Mamba (Tang et al., 2025) and ChoreoMuse (Wang et al., 2025a) predict 2D keypoints from music and then drives image animation with these keypoints. However, they remain limited in handling limb occlusions and complex full-body locomotion in dance videos. In conclusion, existing works for dance video generation still fails to capture the inherently 3D nature of dance, resulting in compromised motion quality and visual appearance. Thus, we propose MACE-Dance, a cascaded expert framework that synergistically integrates motion and appearance generation, producing kinematically plausible and artistically expressive motion while maintaining spatiotemporally coherent visual appearance.
3. Methodology
3.1. Overview
Given a music $M\in R^{T\times C_{m}}$ and reference image $I\in R^{H\times W\times 3}$, our objective is to synthesize the corresponding dance videos $D\in R^{T\times H\times W\times 3}$ with high-quality visual appearance and human motion. Overall, MACE-Dance is with cascaded mixture-of-experts (MoE), as shown in Fig.  2. The Motion Expert (ME) transfers music sequence $M$ into 3D motion sequence $X\in R^{T\times C_{x}}$, enforcing kinematic plausibility and artistic expressiveness. The Appearance Expert (AE) utilizes the above 3D motion sequence $X$ and reference image $I$ to drive video synthesis, preserving visual identity with spatiotemporal coherence. This task decoupling significantly reduces the complexity of learning a direct music-to-video mapping by isolating motion semantics from visual appearance. Moreover, the explicit 3D motion representation suppresses spurious cross-modal correlations and provides an interpretable intermediate interface for robust and controllable video synthesis.
Unlike prior works (Chen et al., 2025e; Tang et al., 2025) that adopt 2D keypoints as the intermediate representation, we instead use 3D motion as the bridge between the two experts for three reasons. (1) Richer spatial fidelity. 3D motion preserves full-body geometric structure, including global translation and orientation, which is essential for dance phrases with large-amplitude locomotion and complex spatial choreography, whereas 2D projections inevitably discard depth and global movement information. (2) Cleaner supervision. 3D representation disentangles pose from camera viewpoint and subject-specific appearance, providing a more stable and generalizable signal for learning the music-to-motion correspondence, while 2D keypoints are entangled with perspective and body proportions. (3) Better robustness. 3D motion is inherently more robust to self-occlusion and viewpoint variation, whereas 2D poses often suffer from missing joints, depth ambiguity, and inconsistent observations. Additionally, we adopt SMPL (Loper et al., 2023) as the representation of the 3D motion sequence $X$ for two reasons. (1) Prior focus on body motion. Most existing 3D dance generation methods primarily model body-level motion rather than detailed hand articulation. In our setting, body-level motion alone is sufficient to produce strong visual results, as also evidenced by our demo videos. (2) Extensibility. Our framework can be readily extended to richer motion representations, such as SMPL-X, when suitable data become available.
Figure 2. Overview of MACE-Dance.Leveraging the cascaded Mixture-of-Experts (MoE) design, the Motion Expert generates kinematically plausible and artistically expressive 3D motion $X$ conditioned on the music $M$, and the Appearance Expert then animates the reference image II with the 3D motion $X$, yielding the dance video $D$ that exhibits spatiotemporally coherent appearance. Thanks to the Guidance-Free Training (GFT) strategy, $\beta\in[0,1]$ can serve as a controllable knob that governs the diversity of the generated motion.
3.2. Motion Expert
3.2.1. Generative Strategy.
Diffusion.
DDPM (Ho et al., 2020) defines diffusion as a Markov noising process with latents $\{z_{t}\}_{t=0}^{T}$ that follow a forward noising process $q(z_{t}|x)$, where $x\sim p(x)$ is drawn from the 3D dance data distribution. The forward noising process is defined as:
(1) $q(z_{t}|x)\sim\mathcal{N}(\sqrt{\bar{\alpha}_{t}}x,(1-\bar{\alpha}_{t})I),$
where $\bar{\alpha}_{t}\in(0,1)$ are constants which follow a monotonically decreasing schedule such that when $\bar{\alpha}_{t}$ approaches 0. Timestep $T$ are commonly set to 1000, and $z_{T}\sim\mathcal{N}(0,I)$. With paired music conditioning $c$, we can reverse the forward diffusion process by learning to estimate $\hat{x}_{\theta}(z_{t},t,c)\approx x$ with model parameters $\theta$ for all $t$. We can optimize $\theta$ by the naive reconstruction loss in Diffusion Model (Ho et al., 2020):
(2) $\mathcal{L}_{\text{DM}}=\mathbb{E}\big[\|\hat{x}_{\theta}(z_{t},t,c)-x\|_{2}^{2}\big].$
Guidance-Free Training.
Conventional classifier-free guidance (CFG (Ho and Salimans, 2022)) modifies the sampling distribution only at inference time by combining conditional and unconditional predictions, which can introduce distribution mismatch and insufficient optimization toward the guided target distribution. In contrast, Guidance-Free Training (GFT (Chen et al., 2025c)) retains the same maximum-likelihood training objective as CFG but adopts a different parameterization that enables a single model to implicitly represent temperature-controlled sampling behavior during training, thereby mitigating distribution mismatch and yielding more stable and consistent high-fidelity generation. Accordingly, we establish $x_{\beta}$ as the new optimization target for our model $\theta$:
(3) $x_{\beta}=\beta\hat{x}_{\theta}(z_{t},t,c,\beta)+(1-\beta)\mathbf{sg}[\hat{x}_{\theta}(z_{t},t,\emptyset,1)]$
where $\emptyset$ denotes the unconditional setting, and $\mathbf{sg}$ represents the stop-gradient operation. $\beta$ serves as a temperature parameter that is also provided to the model $\theta$ as an additional conditioning input. During training, $\beta$ and $t$ are sampled randomly from $U(0,1)$ and the integer set $\{0,1,\dots,T\}$, respectively. Moreover, we further apply the reconstruction loss, 3D joint loss, velocity loss, foot contact loss, to enhance physical plausibility and aesthetic expressiveness:
(4) $\mathcal{L}_{\text{rec}}$ $=\mathbb{E}\big[\|x_{\beta}-x\|_{2}^{2}\big],$
$\mathcal{L}_{\text{joint}}$ $=\mathbb{E}\big[\|FK(x_{\beta})-FK(x)\|_{2}^{2}\big],$
$\mathcal{L}_{\text{vel}}$ $=\mathbb{E}\big[\|FK(x_{\beta})^{\prime}-FK(x)^{\prime}\|_{2}^{2}\big]$
$\mathcal{L}_{\text{foot}}$ $=\mathbb{E}\big[\|FK(x_{\beta})^{\prime}\cdot\hat{\mathbf{b}}\|_{2}^{2}\big],$
where $FK(\cdot)$ denotes the forward kinematic function that converts joint angles into joint positions, and $\hat{\mathbf{b}}$ is the model’s own prediction of the binary foot contact label’s portion of the pose. Our overall training loss $\mathcal{L}$ is the weighted sum of the above losses, where the weights $\lambda$ were chosen to balance the magnitudes of the losses:
(5) $\mathcal{L}=\lambda_{\text{rec}}\mathcal{L}_{\text{rec}}+\lambda_{\text{joint}}\mathcal{L}_{\text{joint}}+\lambda_{\text{vel}}\mathcal{L}_{\text{vel}}+\lambda_{\text{foot}}\mathcal{L}_{\text{foot}}.$
Inference.
At each of the denoising timesteps $t$, Motion Expert predicts the denoised sample and noises it back to timestep t1t-1: $\hat{z}_{t-1}\sim q(\hat{x}_{\theta}(z_{t},t,c,\beta),t-1)$, terminating when it reaches $t=0$. We utilize Denoising Diffusion Implicit Models (DDIM (Song et al., 2021)) to accelerate the sampling procedure. Values of $\beta$ near 0 favor high fidelity, while values near 1 favor high diversity. Thus, $\beta$ can also be regarded as a control signal, and we set its value to 0.75. Notably, GFT inherently achieves theoretically double the generation efficiency compared to conventional CFG, as it only requires a single conditional computation per step, eliminating the need for simultaneous conditional and unconditional predictions.
3.2.2. Model Architecture
Overview.
Motion Expert adopts a BiMamba–Transformer hybrid backbone, thereby enabling the generation of temporally coherent and musically aligned dance motions. BiMamba captures intra-modal local dependencies in music or dance, while the Transformer models cross-modal global context. As shown in Fig.  2, the architecture details are as follows: Firstly, our model conditions the generator on the Librosa (McFee et al., 2015)-extracted music features from $M$ as (Li et al., 2021), which are then processed by an $L_{m}$‑layer BiMamba to capture intra‑modal temporal dynamics. Secondly, the diffusion time step $t$ and temperature parameter $\beta$ are encoded as sinusoidal embeddings and fused by element-wise addition to yield a $t$-$\beta$ embedding used throughout the generator. Third, the dance generator consists of $L_{d}$ stacked blocks. In each block: (1) the current state $z_{t}$ is first passed through a BiMamba to model intra-modal local dependencies; (2) FiLM  (Perez et al., 2018) is applied to modulate the features with the fused $t$-$\beta$ embedding; (3) a Transformer performs cross-modal attention over the music encoding to integrate global musical context, and subsequently passes the result through a feed-forward network; and (4) a second FiLM (Perez et al., 2018) further reinforces the $t$-$\beta$ conditioning. Finally, the generator outputs the 3D motion sequence $\hat{x}_{\theta}(z_{t},t,c,\beta)$ (i.e. $X$ in Sec. 3.1 Overview), represented as SMPL (Loper et al., 2023) parameters. Owing to this architecture, the Motion Expert generates the entire sequence in a non-autoregressive manner during inference, not only improving generation efficiency but also avoiding the exposure-bias problem in autoregressive (Yang et al., 2025b) and inpainting-based (Tseng et al., 2023) methods.
Intra-Modal Local-Dependency.
While the Transformer excels at temporal modeling, it is inherently position-invariant and captures sequence order only through positional encodings (Vaswani, 2017), which limits its deep understanding of local dependencies. In contrast, music-to-dance generation demands strong local continuity between movements. Owing to its inherent sequential inductive bias, Mamba (Gu and Dao, 2023) has demonstrated strong performance in modeling fine-grained local dependencies  (Xu et al., 2024b; Fu et al., 2024). Moreover, Bidirectional Mamba processes inputs in both forward and backward directions, enabling wider representations and deeper understanding of music and dance. Specifically, the Selective State Space Model (Mamba) integrates a selection mechanism and a scan module (S6) (Gu and Dao, 2023) to dynamically emphasize salient input segments for efficient sequence modeling. Unlike traditional SSMs with time-invariant parameters, Mamba generates input-dependent $\bar{A}_{t},\bar{B}_{t},C_{t}$ through fully connected layers, enhancing generalization. For each time step $t$, the input $x_{t}$, hidden state $h_{t}$, and output $y_{t}$ evolve as:
(6) $h_{t}=\bar{A}_{t}h_{t-1}+\bar{B}_{t}x_{t},\quad y_{t}=C_{t}h_{t},$
where $\bar{A}_{t},\bar{B}_{t},C_{t}$ are dynamically updated, and the state transitions become:
(7) $\bar{A}=\exp(\Delta A),\quad\bar{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B,$
where $\Delta$ is the discretization step size, $A$ is the continuous-time state transition matrix, $B$ is the input projection matrix, and $C$ is the output projection matrix.
Cross-Modal Global-Context.
While BiMamba (Gu and Dao, 2023) excels at capturing local dependencies, it is less effective at modeling cross-modal global interactions. Thus, we employ a Transformer (Vaswani, 2017) module after BiMamba in each denoising block, which is crucial for aligning the overall dance structure with long-term musical phrasing. This block consists of a cross-attention layer followed by a feed-forward network (FFN). In the cross-attention layer, motion features serve as queries, while music features provide keys and values:
(8) $\text{Attention}=\text{softmax}\left(\frac{Q_{d}\cdot K_{m}^{T}}{\sqrt{C}}\right)V_{m}.$
In this way, the two components play complementary roles: BiMamba stabilizes short-range intra-modal dynamics, while the Transformer injects cross-modal global musical context to align the generated motion with the overall rhythm and phrase structure.
Figure 3. Qualitative comparison with SOTAs across reference image domains (real-person vs. anime-character) and music genres (Eastern Folk vs. Popping) in the music-driven dance video generation task.
3.3. Appearance Expert
Wan-Animate (Cheng et al., 2025) has recently garnered substantial attention in both industry and academia. However, it is designed for general-purpose motion synthesis; direct transfer to dance video generation is suboptimal due to the domain gap and the richer spatiotemporal complexity of dance—namely intricate whole-body coordination and dynamic camera choreography. Accordingly, the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy to achieve high-fidelity appearance synthesis for dance videos. Specifically, the Kinematic Stage fine-tunes only the Body Adapter while freezing the remaining components, whereas the Aesthetic Stage fine-tunes only the LoRA parameters while keeping the rest of the network fixed.
Model Architecture.
As illustrated in Fig.  2, the architecture of our Appearance Expert is built upon the Wan-Animate (Cheng et al., 2025), which takes a reference image $I$ for appearance and a 3D motion sequence $X$ for motion guidance. The motion sequence $X$ is first projected to 2D keypoints, which are then encoded by a Body Adapter to yield motion features. These features are subsequently fused with the latent extracted from the reference image $I$. The resulting latent is processed by a backbone of stacked DiT blocks, where lightweight LoRA adapters are integrated into each block. The facial processing pipeline remains identical to that of Wan-Animate and is therefore omitted for clarity.
Projector.
We introduce a 3D-to-2D Motion Projector to convert the SMPL sequence generated by the Motion Expert into the 2D pose format required by Wan-Animate. For each frame, we first transform the SMPL parameters into a 3D mesh and render it with pyrender under a fixed frontal-view camera, then apply ViTPose (Xu et al., 2022) to extract the corresponding 2D keypoints. In this way, the projector preserves the benefits of 3D motion modeling while enabling seamless integration with the downstream Appearance Expert.
Kinematic Stage.
In dance, body pose is paramount. The original Wan‑Animate prioritizes facial cues, allocating a dedicated cross‑attention branch to the face while fusing body signals only via additive injection. We therefore strengthen kinematic conditioning by fine‑tuning the Body Adapter in the Kinematic Stage to reweight and calibrate body features across scales, thereby enforcing motion adherence without altering the backbone. We intentionally do not introduce an additional body cross‑attention branch because (i) it disturbs the pretrained inductive bias and can compete with the facial cross‑attention, causing feature entanglement, (ii) it adds substantial memory/latency overhead and training instability on long, fast dance sequences.
Aesthetic Stage.
To refine visual quality without disturbing motion control, we freeze the kinematic pathways and attach lightweight LoRA adapters to the attention (query/key/value/output) and feed-forward projections in each DiT block of Wan‑Animate. These rank‑r adapters enable parameter‑efficient specialization toward dance‑specific aesthetics—sharpening textures (skin, hair, fabric), stabilizing clothing and accessories, and handling rich camera choreography (pans, zooms, handheld motion)—while preserving pretrained content priors. Specifically, LoRA is an effective technique for adapting large pre-trained models to down-streaming tasks with few training-able parameters. To achieve this goal, LoRA introduces a low-rank decomposition-based method to the model’s weight matrix, enabling efficient adaptation to new tasks while maintaining the model’s original capabilities. Given the weight matrix $W_{0}\in\mathbb{R}^{m\times n}$ of the original pre-trained model, LoRA (Hu et al., 2022) uses two low-rank matrices $A\in\mathbb{R}^{m\times r}$ ($r\ll m$) and $B\in\mathbb{R}^{r\times n}$ ($r\ll n$) to shift the trained distribution according to the new data training. Thanks to the low-rank matrix in $A$ and $B$, LoRA updates the model more efficiently than the full rank matrix and shows comparable results with full training. Formally, the new weight matrix $W$ can be represented as:
(9) $W=W_{0}+\Delta W=W_{0}+AB.$
Table 1. Quantitative comparison with SOTAs on the MA-Data dataset in Music-Driven Dance Video Generation task.
Appearance Motion
[…]
Ground Truth 67.12 53.51 91.86 92.97 98.20 96.88 9.24 5.31 0.526
Hallo2 (Cui et al., 2024) [ICLR’25] 62.64 50.79 92.48 93.84 98.30 96.56 16.55 1.29 8.11 5.47 0.505
WAN-S2V (Gao et al., 2025) [Arxiv’25] 64.10 50.20 92.30 93.40 98.20 96.70 18.90 1.45 7.60 5.44 0.485
Echomimic-V3 (Meng et al., 2025) [AAAI’26] 63.20 49.00 91.90 93.10 98.05 96.40 19.60 1.32 7.20 4.60 0.460
EDGE (Li et al., 2023) [CVPR’23] 63.05 49.70 91.79 93.30 98.64 97.10 21.77 1.39 9.08 5.74 0.498
Lodge (Li et al., 2024c) [CVPR’24] 63.69 49.22 91.67 92.98 98.46 97.05 18.73 1.49 8.87 5.71 0.474
MEGA (Yang et al., 2025b) [NeurIPS’25] 66.14 49.89 92.95 94.13 97.45 96.32 18.98 1.65 8.78 5.59 0.513
MACE-Dance 65.35 51.79 93.97 94.57 98.46 97.10 16.46 0.28 9.74 6.34 0.523
Table 2. Quantitative comparison with SOTAs on the FineDance dataset in Music-Driven 3D Dance Generation task.
[…]
Ground Truth 0.216 9.94 7.54 0.201
FACT (Li et al., 2021) [ICCV’21] 113.38 97.05 0.284 3.36 6.37 0.183 29
MNET (Kim et al., 2022) [CVPR’22] 104.71 90.31 0.394 3.12 6.14 0.186 26
Bailando (Siyao et al., 2022) [CVPR’22] 82.81 28.17 0.188 7.74 6.25 0.202 188
EDGE (Tseng et al., 2023) [CVPR’23] 94.34 50.38 0.200 8.13 6.45 0.212 119
Lodge (Li et al., 2024c) [CVPR’24] 50.00 35.52 0.028 5.67 4.96 0.226 224
MEGA (Yang et al., 2025b) [NeurIPS’25] 50.00 13.02 0.243 6.23 6.27 0.226 238
GFT $\rightarrow$ CFG 25.54 36.31 0.295 11.57 7.03 0.223 475
BiMamba $\rightarrow$ Mamba 65.10 51.74 0.35 9.59 7.66 0.224 1044
BiMamba $\rightarrow$ Transformer 104.93 114.42 0.12 7.90 4.22 0.266 683
Motion Expert (Full) 17.83 25.09 0.210 10.30 8.09 0.229 770
4. Experiment
4.1. Dataset
To support music-driven dance video generation task, we curate a large-scale dance video dataset, named MA-Data. It comprises  70k clips of 5–10 seconds each (totaling 116 hours), and spans over 20 distinct dance genres, such as Jazz, Latin, Eastern Folk. MA-Data consists of two complementary sources. (1) 3D-rendered data (motion-centric). This subset is derived from FineDance (Li et al., 2023), the largest 3D dance dataset recorded by professional dancers, and emphasizes professional dance motion rather than visual appearance. Specifically, we first retarget the motion sequence to a character model, then render front-view videos from the 3D character, and extract random 5–10 s segments with a sliding-window strategy for data augmentation, yielding  20k clips ( 28 hours). (2) In-the-wild internet data (appearance-centric). This subset is curated from high-engagement creators on platforms such as TikTok and YouTube, emphasizing visual appearance; motions typically prioritize entertainment value over technical rigor. As raw crawls include many samples misaligned with our task, we apply a multi-stage cleaning pipeline: (i) perform shot boundary detection with TransNet V2 (Soucek and Lokoc, 2024), segment accordingly, and discard segments shorter than 5 s; (ii) remove near-static clips using an optical-flow magnitude threshold; (iii) enforce a single-performer constraint via ViTPose (Xu et al., 2022) by discarding clips that contain multiple people or exhibit little to no human motion; and (iv) split long videos into 5–10 s clips with a sliding window and random offsets. The final set comprises  50k clips ( 88 hours). Finally, we collect an additional 200 5-second clips to construct the test set, with high engagement on TikTok and across multiple dance genres.
4.2. Evaluation
The key challenges of music-driven dance video generation are (Yang et al., 2024c; Zhang et al., 2025a): (1) generating dance motions that are kinematically plausible while artistically expressive; and (2) achieving high-fidelity visual appearance with strong spatiotemporal consistency. Inspired by this, we introduce a motion–appearance evaluation protocol. (1) Motion dimension. We extract 2D keypoint sequences using ViTPose (Xu et al., 2022) from the dance videos and evaluate from a Human-Kinematics perspective. To evaluate the fidelity and diversity, we report FID and DIV across two feature spaces (Li et al., 2021; Siyao et al., 2022): (1) kinetic (k), capturing motion dynamics, and (2) geometric (g), encoding spatial joint relations. To measure music–motion synchronization, we utilize the Beat Alignment Score (BAS) (Li et al., 2021, 2024c). (2) Appearance dimension. Inspired by  (Chen et al., 2025d; Li et al., 2025; Ling et al., 2025), we adopt VBench (Huang et al., 2024)—a widely used benchmark in video generation—and select a set of dance-specific metrics. Our evaluation includes imaging quality (IQ), aesthetic quality (AQ), subject consistency (SC), background consistency (BC), motion smoothness (MS), temporal flickering (TF).
Table 3. Quantitative comparison with SOTAs on the MA-Data dataset in Pose-Driven Image Animation task.
[…]
Animate-Anyone (Hu, 2024) [CVPR’24] 515.26 0.648 0.091 19.65
Magic-Animate (Xu et al., 2024a) [CVPR’24] 1032.06 0.311 0.207 14.00
Wan-Animate (Cheng et al., 2025) [Arxiv’25] 332.82 0.707 0.078 21.11
w/o. Kinematic stage 328.91 0.596 0.107 18.69
w/o. Aesthetic stage 445.93 0.563 0.121 17.89
Appearance Expert 274.94 0.739 0.066 22.40
Figure 4. MACE-Dance generates high-quality dance videos across diverse dance genres.
4.3. Comparison
4.3.1. Music-Driven Dance Video Generation.
As there is currently no open-source implementation for Music-Driven Dance Video Generation, we compare MACE-Dance against two baseline families on the MA-Data dataset: (1) 3D dance generation methods pipelined with Wan-Animate, including EDGE (Tseng et al., 2023), Lodge (Li et al., 2024c), and MEGA (Yang et al., 2025b); (2) General human-motion video generation methods. We perform inference using the pretrained EchoMimic-V3 (Meng et al., 2025) and WAN-S2V (Gao et al., 2025) models, which supports general human motion generation. In addition, we adapt the Hallo2 (Cui et al., 2024) by replacing facial masks with full-body masks, and then fine-tune it on the MA-Data dataset. As shown in Tab.  1, our proposed MACE-Dance demonstrates state-of-the-art (SOTA) performance in both appearance and motion quality. Specifically, for the motion aspect, MACE-Dance achieves the best results across all metrics ($FID_{k}=16.46,FID_{g}=0.28,DIV_{k}=9.74,DIV_{g}=6.34,BAS=0.523$); for the appearance aspect, it also attains best performance on most metrics, with scores of IQ=65.35,AQ=51.79,SC=93.97,BC=94.57,MS=98.46,IQ=65.35,AQ=51.79,SC=93.97,BC=94.57,MS=98.46, and $TF=97.10$. By effectively decoupling the task into 3D Dance Generation and Pose-Driven Image Animation, and leveraging the strong performance of the Motion Expert and Appearance Expert on their respective sub-tasks, MACE-Dance delivers exceptional generation quality. Note: User Study can be found in the supplementary material Sec. 2.
4.3.2. Music-Driven 3D Dance Generation.
Music-Driven 3D Dance Generation is a canonical task and determines the motion quality of MACE-Dance. We compare the Motion Expert against FACT (Li et al., 2021), MNET (Kim et al., 2022), Bailando (Siyao et al., 2022), EDGE (Tseng et al., 2023), Lodge (Li et al., 2024c), and MEGA (Yang et al., 2025b) on the FineDance dataset with metrics following  (Li et al., 2024c). As shown in Tab.  2, the Motion Expert attains overall state-of-the-art (SOTA) performance. Specifically, it achieves the best $FID_{k}=17.83$ and a competitive $FID_{g}=25.09$, indicating high fidelity; the best $DIV_{k}=10.30$ and $DIV_{g}=8.09$, indicating strong diversity; a competitive FSR, supporting physical plausibility; the best $BAS=0.229$, demonstrating superior audio-motion synchronization; and a substantially higher $FPS=770$, evidencing excellent generation efficiency. These advances primarily stem from: (1) adopting a Diffusion Model with a BiMamba-Transformer hybrid architecture, enabling high-quality long motion sequences in a non-autoregressive manner; and (2) a Guidance-Free Training (GFT) strategy that improves generation quality without requiring dual-pass inference (conditioned + unconditioned). Note: Qualitative Comparison can be found in the supplementary material Sec. 3.2.
4.3.3. Pose-Driven Image Animation.
Pose-Driven Image Animation is likewise a canonical task and governs the appearance quality of MACE-Dance. We compare the Appearance Expert against Animate-Anyone (Hu, 2024), Magic-Animate (Xu et al., 2024a), and Wan-Animate (Cheng et al., 2025) on the MA-Data dataset with metrics following  (Cheng et al., 2025). The Appearance Expert achieves state-of-the-art (SOTA) performance on all metrics (Tab. 3, $FVD=274.94$, $SSIM=0.739$, $LPIPS=0.066$, $PSNR=22.40$). Its strong video synthesis quality primarily benefits from Wan-Animate (Baseline)’s powerful cross-modal understanding and the Kinematic-Aesthetic decoupled fine-tuning strategy. Note: Qualitative Comparison can be found in the supplementary material Sec. 3.3.
Figure 5. Ablation for the model architecture of Motion Expert.
Figure 6. MACE-Dance produces coherent long-sequence dance videos.
Figure 7. Ablation for the Appearance Expert.
4.4. Qualitative Analysis
4.4.1. Effect Comparison.
We also present a qualitative comparison against other methods across reference-image domains (real person vs. anime character) and music genres (elegant and rhythmically rich Eastern Folk vs. powerful and funk-inspired Popping) for the music-driven dance video generation task. As shown in Fig.  3, Hallo2 exhibits significant blurring in human details and noticeable artifacts; EDGE often shows abrupt motion discontinuities; Lodge frequently produces abnormal movements that violate physical plausibility; MEGA, WAN-S2V, and Echomimic-V3 often produce overly simple and repetitive motions, limiting expressiveness. In contrast, videos generated by MACE-Dance not only present kinematically plausible and artistically expressive human motion, but also maintain spatiotemporally coherent visual appearance.
4.4.2. Cross-Genre Generation.
Moreover, MACE-Dance generalizes effectively across dance genres as shown in Fig.  4, producing distinct genre-specific motion signatures, including (1) Uyghur dance exhibits light, continuous upper-body rotations with expressive arm trajectories; (2) Dunhuang motion features stable lower-body stances and elegant, circular arm patterns; (3) Dai style emphasizes soft, flowing wrist and elbow movements; (4) K-Pop example demonstrates crisp transitions, symmetrical poses, and rhythm-driven gestures; (5) Popping is characterized by sharp isolations and staccato movements, reflecting its percussive movement vocabulary.
4.4.3. Long-Sequence Generation.
Additionally, a complete music track typically lasts 30 seconds to 5 minutes, making long-sequence generation crucial for practical dance video synthesis (Feng et al., 2025). To mitigate motion drift or visual degradation in long-sequence generation, MACE-Dance incorporates dedicated designs in both stages: (1) a BiMamba–Transformer hybrid in the Motion Expert for drift-free long motion synthesis, and (2) pose-driven relay rendering with identity anchoring in the Appearance Expert. As shown in Fig.  14, MACE-Dance produces coherent long-sequence dance video.
4.5. Ablation Study
4.5.1. Motion Expert
[…]
4.5.2. Appearance Expert
Since the two-stage fine-tuning strategy is central to our Appearance Expert, we evaluate the contribution of each stage via ablation, as summarized in Tab.  3 and Fig.  7. (1) Kinematic Stage. We fine-tune the Body Adapter while freezing all other components. Removing this stage leads to a modest decline across all metrics quantitatively and noticeable kinematic errors and motion blur qualitatively, indicating its effectiveness for ensuring human kinematic plausibility in video. (2) Aesthetic Stage. We fine-tune LoRA parameters in each DiT block. Omitting this stage causes a substantial degradation across all metrics quantitatively and obvious ghosting artifact qualitatively, underscoring its critical role to preserve video aesthetic. Finally, Appearance Expert also demonstrates the superior performance over the Wan-Animate (Baseline), which validates the overall effectiveness of the proposed Kinematic-Aesthetic fine-tuning strategy.
4.5.3. Motion Representation (2D vs. 3D)
Most pose-driven image animation methods rely on 2D keypoints, which naturally motivates using 2D poses as the intermediate motion representation. In contrast, MACE-Dance adopts 3D motion as its intermediate representation. To validate this design, we compare 2D and 3D representations at both the motion-generation level and the final video-generation level. Specifically, for the motion level, we train the same Motion Expert on FineDance with either 2D or 3D motion targets. For the video level, we further render final videos on MA-Data using either 2D pose sequences directly or 3D motion projected to 2D via our projector. As shown in Tab. 4, the 3D-based representation consistently outperforms the 2D-based one across both settings. On FineDance, 3D achieves better fidelity, diversity, and synchronization, indicating that it provides more stable and physically consistent supervision for music-to-motion learning. On MA-Data, 3D further yields clearly better subject consistency, visual fidelity, and beat alignment in the final rendered videos, showing that the advantages of 3D are preserved after the downstream animation stage. These results demonstrate that 3D motion serves as a more reliable intermediate interface than 2D pose for both controllable dance generation and high-quality video synthesis.
Table 4. Comparison of 2D and 3D motion representations.
Kinematic Appearance
[…]
2D 22.8 8.6 6.12 5.24 0.527 51.86 91.84 23.73 0.496
3D 19.5 4.1 8.87 5.92 0.543 51.79 93.97 16.46 0.523
Table 5. Cross-composition analysis of the two experts on MA-Data.
[…]
w/o.ME 50.21 92.10 20.84 0.499
w/o.AE 50.36 91.42 17.92 0.519
Ours 51.79 93.97 16.46 0.523
4.5.4. Role of Each Expert
To further analyze the role of each expert in MACE-Dance, we conduct an additional cross-composition study on MA-Data by replacing one expert at a time with its corresponding baseline counterpart. Specifically, w/o.ME denotes the variant that uses the baseline Motion Expert (EDGE (Tseng et al., 2023)) together with our Appearance Expert, while w/o.AE denotes the variant that uses our Motion Expert together with the baseline Appearance Expert (WAN-Animate (Cheng et al., 2025)). As shown in Tab. 5, the full MACE-Dance consistently achieves the best performance across all evaluated metrics, confirming that both experts contribute positively to the final music-driven dance video generation quality. More specifically, replacing our Appearance Expert with the baseline model leads to clear degradation in appearance-related metrics such as AQ and SC, while replacing our Motion Expert results in a more noticeable drop in the motion-related metric BAS. These observations suggest that the two experts play complementary roles: the Motion Expert mainly strengthens music-motion alignment and body dynamics, whereas the Appearance Expert further improves visual quality, temporal coherence, and identity consistency in the rendered videos.
4.6. Comparison with Video Foundation Models
We further compare MACE-Dance with general-purpose video foundation models, including CogVideoX1.5-5B (Yang et al., 2024d) and WAN2.2-5B (Wan et al., 2025), to examine how a structured motion-to-appearance pipeline performs against recent large-scale video generation models. Although these models demonstrate strong generation ability in broad video domains, they are not specifically designed for music-driven dance video generation, where accurate modeling of beat, rhythm, and body-motion coherence is particularly important. As shown in Table 6, MACE-Dance achieves the best overall performance on SC, FID, and BAS, indicating stronger music-motion alignment, better visual quality, and more consistent identity preservation. Although WAN2.2-5B attains a slightly higher AQ score, it underperforms our method on the other three metrics. Qualitatively, CogVideoX1.5-5B tends to produce weaker and slower dance motions with noticeable blur, while WAN2.2-5B generates larger motion amplitudes but often suffers from temporal identity inconsistency, as shown in Fig. 8. Overall, these results support the effectiveness of explicitly decomposing music-driven dance video generation into a Motion Expert and an Appearance Expert.
Table 6. Comparison with general-purpose video foundation models.
[…]
CogVideoX1.5-5B 50.38 89.92 22.47 0.477
WAN2.2-5B 53.22 90.77 17.53 0.452
Ours 51.79 93.97 16.46 0.523
Figure 8. Comparison with general-purpose video foundation models.
5. Conclusion
In conclusion, we present MACE-Dance, a music-driven dance video generation framework with cascaded Mixture-of-Experts (MoE). The Motion Expert enforces kinematic plausibility and artistic expressiveness, while the Appearance Expert preserves visual identity with spatiotemporal coherence. Specifically, the Motion Expert adopts Diffusion Model with a BiMamba–Transformer hybrid backbone and Guidance-Free Training strategy, while the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy. To better benchmark this task, we curate a large-scale dataset, and design a motion–appearance evaluation protocol. Extensive experiments demonstrate the superiority of MACE-Dance and of its Motion and Appearance Experts. For future work, we plan to extend MACE-Dance with textual descriptions to enable more interactive and flexible dance generation, and improve system-level efficiency to support low-latency authoring and real-time user feedback.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grants 62436010, 72572090, 62572474 and 62172421, and in part by Tsinghua University School of Economics and Management Research Grant.